# Excluding cosmetic surgeries on hyperbolic 3-manifolds

### v2

@@ 6.1. Arithmetic definitions
A number field, denoted $k$, is a finite extension of ${\mathbb{Q}}$. The ring of integers of a number field $k$ is denoted $R_{k}$. A quaternion algebra over $k$ is a $4$–dimensional $k$–vector space with basis $1,i,j,ij$, with multiplication defined so that $i^{2}=a\in k$ and $j^{2}=b\in k$ and $ij=-ji$.
A place of a number field $k$ is an equivalence class of valuations $v:k\to{\mathbb{R}}$. A place is called real if it arises from the absolute value ||$\cdot$|| ⋅ | in an embedding $k\hookrightarrow{\mathbb{R}}$, meaning $v(x)=|x|$. A place is called complex if it arises from the absolute value ||$\cdot$|| ⋅ | in a non-real embedding $k\hookrightarrow{\mathbb{C}}$. A place $v$ is called finite or non-Archimedean if it satisfies $v(x+y)\leq\max(v(x),v(y))$ for all $x,y\in k$. Every place is real, complex, or finite, and finite places correspond to prime ideals $\mathcal{P}\subset R_{k}$ [36, Theorem 0.6.6].
A quaternion algebra $A$ over $k$ is ramified at finitely many places of $k$ [36, Definitions 2.7.1]. In fact, $A$ is determined by the set of places over which it is ramified [36, Theorem 2.7.5]. The finite ramification locus $\text{Ram}_{f}(A)$ is the set prime ideals $\mathcal{P}\subset R_{k}$ corresponding to the finite places where $A$ is ramified.
@@
The invariant trace field of $\Gamma$ is $k(\Gamma)={\mathbb{Q}}(\operatorname{tr}\Gamma^{(2)})$. This is always a number field [36, Theorem 3.1.2]. In addition, $k(\Gamma)$ is an invariant of $\mathcal{C}(\Gamma)$ [36, Theorem 3.3.4].
Let $g,h\in\Gamma^{(2)}$ be any pair of non-commuting loxodromic elements. We lift $g,h\in PSL(2,{\mathbb{C}})$ to matrices in $SL(2,{\mathbb{C}})$. The invariant quaternion algebra $A(\Gamma)$ is the $4$–dimensional algebra over $k(\Gamma)$ with basis $I,g,h,gh$:
@@
That this is a quaternion algebra is established in [36, Corollary 3.2.3]. While the choice of basis elements $g,h$ is highly non-unique, the algebra $A(\Gamma)$ is also an invariant of the commensurability class $\mathcal{C}(\Gamma)$ [36, Corollary 3.3.5]. When $\Gamma$ is arithmetic (see definition below), the pair $k(\Gamma)$ and $A(\Gamma)$ form a complete invariant of the commensurability class $\mathcal{C}(\Gamma)$.
@@
The equivalence between arithmeticity and these conditions is [36, Theorem 8.3.2].
It is also possible to produce a Kleinian group directly from a number field $k$ and a quaternion algebra $A$ over $k$ satisfying conditions $(1)$$(2)$. An order in $A$ is a ring of integers $\mathcal{O}\subset A$ that contains $R_{k}$ and satisfies $k\mathcal{O}=A$. An order is maximal if it is maximal with respect to inclusion. The set of units in a maximal order $\mathcal{O}$ is denoted $\mathcal{O}^{1}$. Orders in a quaternion algebra play a similar role to rings of integers in a field. Now, the single non-real embedding $k\hookrightarrow{\mathbb{C}}$ induces an embedding $\rho\colon A\to\operatorname{Mat}(2,{\mathbb{C}})$ into the set of $2\times 2$ matrices over ${\mathbb{C}}$. The image $\rho(A)$ is typically indiscrete, but restricting the domain to $\mathcal{O}^{1}$ produces a discrete, faithful representation $\rho\colon\mathcal{O}^{1}\to SL(2,{\mathbb{C}})$. Projecting down to $PSL(2,{\mathbb{C}})$ produces a Kleinian group derived from a quaternion algebra. This group, denoted $\Gamma_{\mathcal{O}^{1}}=P\rho(\mathcal{O}^{1})$, has integer traces and is therefore arithmetic.
@@ 6.2. Volumes
The following is a special case of a foundational theorem due to Borel [36, Theorem 11.1.3].
@@
Since $k={\mathbb{Q}}(e^{\pi i/3})={\mathbb{Q}}(\sqrt{-3})$ is a quadratic imaginary extension of ${\mathbb{Q}}$, it has no real places and exactly one complex place. Any quaternion algebra $A$ over $k$ is ramified at the empty set of real places, so $k$ and $A$ satisfy conditions $(1)$$(2)$.
To compute $\operatorname{vol}({\mathbb{H}}^{3}/\Gamma_{\mathcal{O}^{1}})$ using Borel’s volume formula, we need the following information. The field $k$ has discriminant $\Delta_{k}=-3$ by [36, Example 0.2.7]. The degree of the extension is $[k:{\mathbb{Q}}]=2$. The Dedekind zeta function $\zeta_{k}(2)$ was computed by Zagier [47, page 291] in terms of the Lobachevsky function $\Lambda$:
@@
With this information, Borel’s volume formula [36, Theorem 11.1.3] gives
@@
We can now prove a result that is needed in Theorems 5.1 and 5.2.
Proposition 6.2.
Let $Q$ be the closed 3-manifold produced by $(8,1)$ filling on the $5_{2}$ knot. Let $R$ be the closed 3-manifold produced by $(56,3)$ Dehn filling on the $(-2,3,7)$ pretzel knot. Then both $Q$ and $R$ are hyperbolic and arithmetic. Furthermore,
$\operatorname{vol}(Q)=\operatorname{vol}(R)=2V_{3}.$
We note that the $5_{2}$ knot appears in the SnapPy census as m015. The $(-2,3,7)$ pretzel appears in the knot tables as 12n242 and in the SnapPy census as m016.
Proof.
Using SnapPy inside Sage, we verify that $Q$ and $R$ are hyperbolic and estimate their volumes to within rigorously bounded error.
@@
Thus $k(\Gamma)={\mathbb{Q}}(\alpha)$ where $\alpha=e^{\pi i/3}$ is a root of $x^{2}-x+1$. This field has exactly one complex place (because the two roots are complex conjugates) and no real places. Thus conditions $(1)$$(2)$ above are satisfied. Snap also checks that $\Gamma$ has integer traces, so $(3)$ is satisfied and $\Gamma$ is arithmetic.
@@
Next, we estimate the smallest covolume of any lattice in $\mathcal{C}(\Gamma)$. By [36, Theorem 11.5.2], the smallest covolume is realized by a unique (up to conjugation) Kleinian group $\Gamma_{\mathcal{O}}=\Gamma_{\emptyset,\mathcal{O}}$.
By [36, Theorem 11.6.3], we learn $\Gamma_{\mathcal{O}^{1}}$ is a normal subgroup of $\Gamma_{\mathcal{O}}$, with quotient an abelian $2$–group. Since the ring of integers of $k(\Gamma)$, namely ${\mathbb{Z}}[e^{\pi i/3}]$, is a principal ideal domain, $k(\Gamma)$ has class number $1$, hence [36, Theorem 11.6.5 and Corollary 11.6.4] give
$[\Gamma_{\mathcal{O}}:\Gamma_{\mathcal{O}^{1}}]=2^{n}\qquad\text{where}\qquad n\leq 1+|\operatorname{Ram}_{f}(A)|=3.$
@@
Finally, observe that the only prime in $k(\Gamma)$ that divides $2$ is $(2)$ itself, and this prime ideal is not ramified in $A(\Gamma)$. Thus [36, Theorem 11.5.2] says that every Kleinian group $\Gamma$ commensurable to $\Gamma_{\mathcal{O}^{1}}$ must have volume an integer multiple of $\frac{1}{2}\operatorname{vol}({\mathbb{H}}^{3}/\Gamma_{\mathcal{O}})$. Combining this with (6.4) gives
@@
Remark 6.5.
There is an alternate way to establish the equality
$\operatorname{vol}(Q)=\operatorname{vol}(R)=2V_{3}.$
[…]
James Clift has written code that can perform a sequence of 22223333 moves, 33332222 moves, and 222200 moves (cancellations), to verify that the three-tetrahedron spun triangulation of $Q$ is scissors congruent to the two-tetrahedron triangulation of the figure–$8$ knot complement. His code uses Snap to follow exact solutions to the gluing equations through the retriangulation moves, with shapes given in the number field $k(\Gamma)={\mathbb{Q}}(e^{\pi i/3})$. The same argument works for $R$. Since the manifolds are scissors congruent, they have equal volumes. The output of Clift’s code is reproduced in the ancillary files [13].
@@ 7. Additional questions
Finally, we observe that in Theorem 5.1, the number of $n$–crossing knots that share a common filling with the figure–$8$ knot does not seem to increase with $n$. By contrast, the total number of $n$–crossing knots grows exponentially. Hence, we wonder:
@@
For $n\geq 4$, let $\mathcal{K}_{n}$ be the set of hyperbolic knots of $n$ crossings that share a common filling with the figure–$8$ knot. By [32, Theorem 1.2], every $\mathcal{K}_{n}$ contains at least one double twist knot, so it is nonempty. Is the cardinality $|\mathcal{K}_{n}|$ universally bounded as $n$ grows? The same question also makes sense if the figure–$8$ knot is replaced by any fixed $K$.

### v3

@@ 6.1. Arithmetic definitions
A number field, denoted $k$, is a finite extension of $\mathbb{Q}$}. The ring of integers of a number field $k$ is denoted $R_{k}$. A quaternion algebra over $k$ is a 44–dimensional kk–vector space with basis $1,i,j,ij$, with multiplication defined so that $i^{2}=a\in k$ and $j^{2}=b\in k$ and ij=jiij=-ji.
A place of a number field $k$ is an equivalence class of valuations $v:k\to{\mathbb{R}}$. A place is called real if it arises from the absolute value $|\cdot|$ in an embedding $k\hookrightarrow{\mathbb{R}}$, meaning $v(x)=|x|$. A place is called complex if it arises from the absolute value $|\cdot|$ in a non-real embedding $k\hookrightarrow{\mathbb{C}}$. A place $v$ is called finite or non-Archimedean if it satisfies $v(x+y)\leq\max(v(x),v(y))$ for all $x,y\in k$. Every place is real, complex, or finite, and finite places correspond to prime ideals $\mathcal{P}\subset R_{k}$ [41, Theorem 0.6.6].
A quaternion algebra $A$ over $k$ is ramified at finitely many places of $k$ [41, Definitions 2.7.1]. In fact, $A$ is determined by the set of places over which it is ramified [41, Theorem 2.7.5]. The finite ramification locus $\text{Ram}_{f}(A)$ is the set of prime ideals $\mathcal{P}\subset R_{k}$ corresponding to the finite places where $A$ is ramified.
@@
The invariant trace field of $\Gamma$ is $k(\Gamma)={\mathbb{Q}}(\operatorname{tr}\Gamma^{(2)})$. This is always a number field [41, Theorem 3.1.2]. In addition, $k(\Gamma)$ is an invariant of $\mathcal{C}(\Gamma)$ [41, Theorem 3.3.4].
Let $g,h\in\Gamma^{(2)}$ be any pair of non-commuting loxodromic elements. We lift $g,h\in PSL(2,{\mathbb{C}})$ to matrices in $SL(2,{\mathbb{C}})$. The invariant quaternion algebra $A(\Gamma)$ is the 44–dimensional algebra over $k(\Gamma)$ with basis $I,g,h,gh$:
@@
That this is a quaternion algebra is established in [41, Corollary 3.2.3]. While the choice of basis elements $g,h$ is highly non-unique, the algebra $A(\Gamma)$ is also an invariant of the commensurability class $\mathcal{C}(\Gamma)$ [41, Corollary 3.3.5]. When $\Gamma$ is arithmetic (see definition below), the pair $k(\Gamma)$ and $A(\Gamma)$ form a complete invariant of the commensurability class $\mathcal{C}(\Gamma)$.
@@
The equivalence between arithmeticity and these conditions is [41, Theorem 8.3.2].
It is also possible to produce a Kleinian group directly from a number field $k$ and a quaternion algebra $A$ over $k$ satisfying conditions (1)(1)$(2)$. An order in $A$ is a ring of integers $\mathcal{O}\subset A$ that contains $R_{k}$ and satisfies $k\mathcal{O}=A$. An order is maximal if it is maximal with respect to inclusion. The set of units in a maximal order $\mathcal{O}$ is denoted $\mathcal{O}^{1}$. Orders in a quaternion algebra play a similar role to rings of integers in a field. Now, the single non-real embedding $k\hookrightarrow{\mathbb{C}}$ induces an embedding $\rho\colon A\to\operatorname{Mat}(2,{\mathbb{C}})$ into the set of $2\times 2$ matrices over $\mathbb{C}$}. The image $\rho(A)$ is typically indiscrete, but restricting the domain to $\mathcal{O}^{1}$ produces a discrete, faithful representation $\rho\colon\mathcal{O}^{1}\to SL(2,{\mathbb{C}})$. Projecting down to $PSL(2,{\mathbb{C}})$ produces a Kleinian group derived from a quaternion algebra. This group, denoted $\Gamma_{\mathcal{O}^{1}}=P\rho(\mathcal{O}^{1})$, has integer traces and is therefore arithmetic.
@@ 6.2. Volumes
The following is a special case of a foundational theorem due to Borel [41, Theorem 11.1.3].
@@
Since $k={\mathbb{Q}}(e^{\pi i/3})={\mathbb{Q}}(\sqrt{-3})$ is a quadratic imaginary extension of $\mathbb{Q}$}, it has no real places and exactly one complex place. Any quaternion algebra $A$ over $k$ is ramified at the empty set of real places, so $k$ and $A$ satisfy conditions (1)(1)$(2)$.
To compute $\operatorname{vol}({\mathbb{H}}^{3}/\Gamma_{\mathcal{O}^{1}})$ using Borel’s volume formula, we need the following information. The field $k$ has discriminant $\Delta_{k}=-3$ by [41, Example 0.2.7]. The degree of the extension is $k:{\mathbb{Q}}]=2$. The Dedekind zeta function $\zeta_{k}(2)$ was computed by Zagier [52, page 291] in terms of the Lobachevsky function $\Lambda$:
@@
With this information, Borel’s volume formula [41, Theorem 11.1.3] gives
@@
We can now prove Proposition 6.1. Recall that the $5_{2}$ knot appears in the SnapPy census as m015. The (2,3,7)(-2,3,7) pretzel appears in the knot tables as 12n242 and in the SnapPy census as m016.
Proof of Proposition 6.1.
Using SnapPy inside Sage, we verify that $Q=\mathtt{5\_2(8,1)}$ and $R=\mathtt{12n242(56,3)}$ are hyperbolic and estimate their volumes to within rigorously bounded error.
@@
Thus $k(\Gamma)={\mathbb{Q}}(\alpha)$ where $\alpha=e^{\pi i/3}$ is a root of $x^{2}-x+1$. This field has exactly one complex place (because the two roots are complex conjugates) and no real places. Thus conditions (1)(1)$(2)$ above are satisfied. Snap also checks that $\Gamma$ has integer traces, so $(3)$ is satisfied and $\Gamma$ is arithmetic.
@@
Next, we estimate the smallest covolume of any lattice in $\mathcal{C}(\Gamma)$. By [41, Theorem 11.5.2], the smallest covolume is realized by a unique (up to conjugation) Kleinian group $\Gamma_{\mathcal{O}}=\Gamma_{\emptyset,\mathcal{O}}$.
By [41, Theorem 11.6.3], we learn $\Gamma_{\mathcal{O}^{1}}$ is a normal subgroup of $\Gamma_{\mathcal{O}}$, with quotient an abelian 22–group. Since the ring of integers of $k(\Gamma)$, namely ${\mathbb{Z}}[e^{\pi i/3}]$, is a principal ideal domain, $k(\Gamma)$ has class number $1$, hence [41, Theorem 11.6.5 and Corollary 11.6.4] give
$\Gamma_{\mathcal{O}}:\Gamma_{\mathcal{O}^{1}}]=2^{n}\qquad\text{where}\qquad n\leq 1+|\operatorname{Ram}_{f}(A)|=3.$
@@
Finally, observe that the only prime in $k(\Gamma)$ that divides $2$ is $(2)$ itself, and this prime ideal is not ramified in $A(\Gamma)$. Thus [41, Theorem 11.5.2] says that every Kleinian group $\Gamma$ commensurable to $\Gamma_{\mathcal{O}^{1}}$ must have volume an integer multiple of $\frac{1}{2}\operatorname{vol}({\mathbb{H}}^{3}/\Gamma_{\mathcal{O}})$. Combining this with (6.5) gives
@@ 7. Additional questions
Finally, we observe that in Theorem 5.1, the number of nn–crossing knots that share a common filling with the figure–$8$ knot does not seem to increase with $n$. By contrast, the total number of nn–crossing knots grows exponentially. Hence, we wonder:
@@
For $n\geq 4$, let $\mathcal{K}_{n}$ be the set of hyperbolic knots of $n$ crossings that share a common filling with the figure–$8$ knot. By [37, Theorem 1.2], every $\mathcal{K}_{n}$ contains at least one double twist knot, so it is nonempty. Is the cardinality $|\mathcal{K}_{n}|$ universally bounded as $n$ grows? The same question also makes sense if the figure–$8$ knot is replaced by any fixed $K$.